# Berezinskii-Kosterlitz-Thouless transition in the XY model on the honeycomb lattice: A comprehensive Monte Carlo analysis

### v2

@@ Abstract
In this paper, we examined the Berezinskii-Kosterlitz-Thouless (BKT) phase transition in the two-dimensional XY model on the honeycomb lattice. To address its thermodynamical behavior, we conducted standard numerical Monte Carlo simulations with the Simulated Annealing (SA) protocol and entropic simulations based on the Wang-Landau (WL) algorithm. The inverse transition temperature ($\beta_{BKT}$) was determined using the second ($\Upsilon$) and fourth-order ($\Upsilon_{4}$) helicity modulus as the order parameter. We also used the fourth-order Binder cumulant of the magnetization as an additional tool to address the critical properties of the model. Our finite-size scaling (FSS) analysis suggests $\beta_{BKT}=1.696(3)$ from $\Upsilon$ and $\beta_{BKT}=1.67(1)$ from $\Upsilon_{4}$. From the Binder cumulant crossing we found $\beta_{BKT}=1.724(2)$. These values deviate significantly from the expected theoretical value of $\sqrt{2}$. We believe that this discrepancy suggests that the theoretical assumptions regarding the analytical calculation may need to be revisited. Additionally, we calculated the vortex density and the formation energy of the vortex-antivortex pairs, where the obtained vortex formation energy is $2\mu=4.0(5)$. Upon comparison with the square lattice, our results support the notion of instability of the honeycomb lattice and give additional backing to the critical behavior we found.
@@ I Introduction
In 2016, scientists David J. Thouless, F. Duncan M. Haldane, and J. Michael Kosterlitz were awarded the Nobel Prize in Physics for their discoveries regarding topological phase transitions in matterNob . They demonstrated that, at sufficiently low temperatures, the magnetic moments of a two-dimensional lattice of atoms form vortex-antivortex pairs because of topological defects. However, a topological phase transition occurs when the system temperature increases. At a certain critical temperature ($\beta_{BKT}$), these pairs dissociate and roam freely through the lattice. This transition is called the Kosterlitz-Thouless (KT) transition or BKT, in honor of Vadim Berezinskii, a deceased theoretical physicist who presented ideas similar to Kosterlitz and ThoulessBerezinskii (1972).
@@
This study is motivated by two key factors. First, the BKT transition temperature for the square lattice is well established in the literature, with numerical calculations yielding $0.8935(1)$Hsieh et al. (2013). For the honeycomb lattice, some analytical studies predicts an exact value of $1/\sqrt{2}\approx 0.707$Nienhuis (1982); Campostrini et al. (1996), which is confirmed by later calculations using cluster simulationsDeng et al. (2007) and diffusion map methodWang et al. (2021). However, it seems that a comprehensive Monte Carlo (MC) analysis of the critical behavior on the honeycomb lattice is missing. The second motivation stems from the ongoing debate regarding the appropriate parameter to characterize the BKT transition, given the limitations of computer simulations due to the finite size of simulated systems. The modulus of second-order helicity, which measures the system response to a global twist of spins in a given direction, is currently accepted as the parameter for describing this transitionMinnhagen and Kim (2003). The controversy arises because the discontinuity in the helicity modulus is not strongly dependent on the system size. While some researchers have shown that logarithmic corrections applied to the helicity modulus can yield accurate estimates of $\beta_{BKT}$sieh et al. (2013), it is also possible to use fourth-order helicity as an order parameter. This alternative presents a discontinuity at the critical temperature, potentially providing a more precise description of the BKT transition via entropic numerical simulations, where a high resolution search for the critical points is possibleCaparica (2014).
We primarily use the MC method along with the Simulated Annealing (SA) protocolLandau and Binder (2014) to calculate the second- and fourth-order helicity. This helps us to estimate the inverse critical temperature $\beta_{BKT}$. Additionally, we use entropic simulations based on the Wang-Landau (WL) algorithmWang and Landau (2001a) to validate the SA results. The WL technique allows for a more precise determination of the critical region without being affected by critical slowdown, although it is limited to small system sizes for models with continuous symmetry. Moreover, we will employ the finite-size scaling (FSS) technique with logarithmic correction as proposed by Y.D. Hsieh et al.Hsieh et al. (2013). We also calculate the vortex density and the energy formation of vortex pairs, which may provide insight into the stability of the honeycomb lattice compared to the square lattice and the location of the critical region.
@@ II The model
The XY model represents a system of spins or vectors arranged at the sites of a two-dimensional latticeKosterlitz and Thouless (1973) and serves as a powerful tool for modeling magnetic systems in two dimensions, such as magnetic crystalsGong and Zhang (2019) and van der Waals heterostructuresGong et al. (2017).
@@
However, in the XY model, there is no discontinuity in the first derivatives of the free energy at the transition temperature, but rather in the second derivatives. Furthermore, there is no spontaneous symmetry breaking of the system, but rather a change in the nature of quasi-long-range fluctuations of the order parameter, i.e., there are no abrupt changes of the system during the transition. This configures a phase transition of continuous (infinite) orderBerezinskii (1972); Kosterlitz (1974); Kosterlitz and Thouless (1973). For sufficiently low temperatures, there is the creation of vortex-antivortex pairs in the lattice, representing a state with almost long-range order, as predicted by Kosterlitz and ThoulessKosterlitz and Thouless (1973). Vortices can be characterized as topological defectsNob , where the spins arrange themselves in a whirlpool-like formation, that is, when adding the phase differences between the spins in a closed path proportional to the lattice geometry, it yields a value multiple of ±$\pm 2\pi$. As the temperature of the system increases, therefore, a transition from this state to a disordered state occurs, breaking the vortex-antivortex pairs, which begin to move freely through the system.
@@ III.1 The Metropolis Algorithm
The Monte Carlo (MC) method uses random numbers to construct the phase space of a given physical system, which can be studied by sampling its microstates. This capability of the MC method is convenient when one wants to represent complex and random systems with fair precisionLandau and Binder (2014).
@@
By scanning the entire lattice, that is, all $N$ sites, where $N=L^{2}$, assigning a new orientation to the spin at each site and using the Metropolis algorithm to accept or not the new orientation, we perform which is known as the Monte Carlo sweep ($m_{s}$).
@@ III.2 Simulated Annealing
Due to the complexity of the phase space of the XY model, the Metropolis algorithm may encounter difficulties in sampling the low temperature region. A technique used to overcome this type of problem is called Simulated Annealing (SA)Landau and Binder (2014). This technique consists of initializing the system at an infinite temperature and cooling it uniformly and slowly until $T\approx 0$. This protocol is efficient because it allows the system not to get stuck in local energy minima when you want to minimize a function dependent on a very large number of parameters, since such functions present difficulties in locating the global minimum due to the high number of local minima. Therefore, the SA procedure allows the system to pass through local minima without getting stuck in them, as long as the increase in $\beta=1/T$ is small in each interval.
In practice, we start with $\beta=0$ and increase it uniformly until we reach a desired temperature value. Each simulation consists of $n$ temperature intervals, where within each interval $m_{s}$ Monte Carlo sweeps are performed, resulting in a total of $nm_{s}$ simulation steps. After each Monte Carlo step, the increments in $\beta$ are given by $\beta=1/k_{B}T=\beta_{0}+i\Delta\beta_{n}\quad(i=1,...,nm_{s})$, where $\Delta\beta_{n}=(\beta_{n}-\beta_{0})/nm_{s}$.
@@ III.3 Wang-Landau simulation
In the study of phase transitions in physical systems, entropic simulations have proven to be an extremely useful toolCaparica and DaSilva (2015); Caparica and Cunha-Netto (2012). Based on the Wang-Landau scheme, they are performed by a random walk in energy space, driven by an unbiased change in spin states, to construct the density of states $g(E)$. The probability of accepting a new spin state is proportional to the reciprocal of the density of states, $1/g(E)$. During the simulation, a histogram of energy values, $H(E)$, is accumulated, and the usual flatness criterion for $H(E)$ is set $90\%$. After estimating the density of states, these simulations allow the calculation of average values of relevant thermodynamic quantities at any temperature, significantly expanding the information available about the system. In our work, we use entropic sampling based on the Wang-Landau methodWang and Landau (2001a, b), with some improvements, such as the adoption of the Monte Carlo sweep to update the density of states, the accumulation of microcanonical average is done only after a more advanced WL level, the use of an automatic criterion to end the simulations, and the use of an improvement that saves computational time, which consists of starting all the rounds from a single previous simulation up to some WL level and performing all the other runs from thereCaparica and Cunha-Netto (2012); Caparica (2014); Caparica and DaSilva (2015).
@@ III.4 Thermodynamic quantities
To describe the thermodynamic behavior of the XY model, we computed the average energy per site, the magnetization, and the corresponding thermal fluctuations, that is, specific heat and susceptibility. We also analyzed the behavior of the fourth-order cumulant of the order parameter. The average energy in this case is given by:
@@
where $N$ is the total number of spins, $\cal H$ is the Hamiltonian of the system given by Equation 1 and its average is taken at a certain temperature.
Although magnetization is not the appropriate order parameter for obtaining the BKT transition temperature, it is important for a qualitative understanding of the phase transition. Thus, we can write the magnetization of the XY model in a finite lattice as:
@@
With this definition, $m=1$ in the ground state and should decrease as the temperature increases. For the specific heat and susceptibility we have,
@@
In turn, it is argued that the appropriate order parameter for studying the BKT transition is the second-order helicity modulusMinnhagen and Kim (2003); Hsieh et al. (2013), which corresponds to a response of the system to a global twist of the spins in a certain direction. To better understand helicity, one can imagine a square lattice of spins at a temperature $T=0$ in the ferromagnetic phase with linear size $L$, and a magnetic field parallel to one of its edges is applied. The spins located at the edge that is under the action of the magnetic field must undergo a change of $\Phi$ in their orientation relative to the original state. To minimize the energy of the system, each spin of neighboring rows parallel to the edge will undergo a rotation equivalent to an angle of $\phi=\Phi/(L-1)$. Thus, the energy of this system in the presence of this spin is given by $E(\phi)=-J(L-1)L\cos(\phi)=E(0)+J(L-1)L[1-\cos(\phi)]$. According to Sandvik et al. (2010), for a small $\phi$, we have $E(\phi)-E(0)=(J/2)(L-1)L\phi^{2}$. Using this result, one can see that the change in the energy for a small twist corresponds to a function of $\phi^{2}$, thus, the helicity for $T=0$ is defined as
@@
because the odd terms of the expansion are null. Thus, we define helicity and fourth-order helicity modulus as follows:
@@
According to Minnhagen and Kim (2003), for a potential that depends only on the difference in the angle between the spins, the helicity is given by:
@@
$s=\frac{1}{N}\sum_{\langle ij\rangle_{x}}\sin(\theta_{i}-\theta_{j})$ (19)
@@
Therefore, $\langle\Upsilon\rangle\geq 0$, since for small variations of $\phi$, the free energy is dominated by the smallest non-zero derivative of $F$, which is positive. However, $\langle\Upsilon_{4}\rangle$ must also be positive for any temperature $T$ where $\langle\Upsilon\rangle$ is equal to zero. This implies that $\langle\Upsilon\rangle$ cannot continually go to zero at the transition temperature $T_{BKT}$ if $\langle\Upsilon_{4}\rangle$ at the same time approaches a minimum at $T_{BKT}$. Therefore, if $\langle\Upsilon_{4}\rangle$ approaches the minimum in $T_{BKT}$, then the change in $\langle\Upsilon\rangle$ must be discontinuousMinnhagen and Kim (2003). Thus, as demonstrated by Minnhagen, both orders of helicity can be used as an order parameter in the BKT transitionMinnhagen and Kim (2003).
@@
This equality is satisfied when the system size goes to infinity. Thus, the procedure is to calculate the helicity for each temperature at different lattice sizes and compare with the curve $r=2/\beta\pi$. The intersection of these values gives us the critical temperature value for each lattice size. To estimate the BKT transition temperature using fourth-order helicity (Equation 20), we simply need to identify the temperature that minimizes the function.
@@
To overcome the limitation of the simulation to the size of the system, we will use the technique known as Finite-Size Scaling (FSS), developed by K. BinderBinder (1981). For temperatures below the critical value, some physical quantities of interest, such as helicity $\Upsilon$, can be described by a power lawSilva et al. (2022), as $\Upsilon\propto L^{-x}$, where $x$ is an exponent dependent on the system temperature. When the order parameter has a strong dependence on the size of the system and it is possible to find a singularity in the phase transition, a linear extrapolation ($x=1$) is appropriate to find the critical temperature in the thermodynamic limit ($L\rightarrow\infty$). Otherwise, it cannot be guaranteed that the order parameter is discontinuous in this limit. This appears to be the case for second-order helicity. However, to try to get around this problem, some authors use a logarithmic correction given byLandau and Binder (2014); Silva et al. (2022); Binder (1997):

### v3

@@ Abstract
In this paper, we examined the Berezinskii-Kosterlitz-Thouless (BKT) phase transition in the two-dimensional XY model on the honeycomb lattice. To address its thermodynamical behavior, we conducted standard numerical Monte Carlo simulations with the simulated annealing protocol and entropic simulations based on the Wang-Landau algorithm. The inverse transition temperature ($\beta_{BKT}$) was determined using the second ($\Upsilon$) and fourth-order ($\Upsilon_{4}$) helicity modulus as the order parameter. We also used the fourth-order Binder cumulant of the magnetization as an additional tool to address the critical properties of the model. Our finite-size scaling analysis suggests $\beta_{BKT}=1.687(3)$ from $\Upsilon$ and $\beta_{BKT}=1.635(11)$ from $\Upsilon_{4}$. From the Binder cumulant crossing we found $\beta_{BKT}=1.724(2)$. These values deviate significantly from the expected theoretical value of $\sqrt{2}$. We believe that this discrepancy suggests that the theoretical assumptions regarding the analytical calculation may need to be revisited. Additionally, we calculated the vortex density and the formation energy of the vortex-antivortex pairs, where the obtained vortex formation energy is $2\mu=4.0(5)$. Upon comparison with the square lattice, our results support the notion of instability of the honeycomb lattice to support the spin long-range order and give additional backing to the critical behavior we found.
@@ I Introduction
In 2016, the Nobel Prize in Physics was awarded to David J. Thouless, F. Duncan M. Haldane, and J. Michael Kosterlitz for their groundbreaking contributions to the understanding of topological phase transitions and phases of matter, which have significantly advanced condensed matter physics. They demonstrated that, at sufficiently low temperatures, the magnetic moments of a two-dimensional lattice of atoms form vortex-antivortex pairs due to the presence of topological defects. However, a topological phase transition occurs when the system temperature increases. At a certain critical temperature ($\beta_{BKT}$), these pairs dissociate and roam freely through the lattice. This transition is called Kosterlitz-Thouless (KT) transition or BKT, in honor of Vadim Berezinskii, a deceased theoretical physicist who presented ideas similar to Kosterlitz and ThoulessBerezinskii (1972).
@@
The present study is motivated by two key factors. First, the BKT transition temperature for the square lattice is well established in the literature, with numerical calculations yielding $0.8935(1)$Hsieh et al. (2013). For the honeycomb lattice, some analytical studies predict an exact value of $1/\sqrt{2}\approx 0.707$Nienhuis (1982); Campostrini et al. (1996), which is confirmed by later calculations using cluster simulationsDeng et al. (2007) and diffusion map methodWang et al. (2021). Nonetheless, a detailed Monte Carlo (MC) study of the critical behavior on the honeycomb lattice appears to be absent in the existing literature. The second motivation stems from the ongoing debate regarding the appropriate parameter to characterize the BKT transition, given the limitations of computer simulations due to the finite size of simulated systems. The modulus of second-order helicity, which measures the system response to a global twist of spins in a given direction, is currently accepted as the parameter to describe this transitionMinnhagen and Kim (2003). The controversy arises because the discontinuity in the helicity modulus is not strongly dependent on the system size. While some researchers have shown that logarithmic corrections applied to the helicity modulus can yield accurate estimates of $\beta_{BKT}$sieh et al. (2013), it is also possible to use fourth-order helicity as an order parameter. This alternative presents a discontinuity at the critical temperature, potentially providing a more precise description of the BKT transition via entropic numerical simulations, where a high resolution search for the critical points is possibleCaparica (2014).
We primarily use the MC method and the Simulated Annealing (SA) protocolLandau and Binder (2014) to calculate the second- and fourth-order helicity. This helps us to estimate the inverse critical temperature $\beta_{BKT}$. Additionally, we use entropic simulations based on the Wang-Landau (WL) algorithmWang and Landau (2001a) to validate the SA results. The WL technique allows for a more precise determination of the critical region without being affected by critical slowdown. However, it is limited to small system sizes for models with continuous symmetry. Moreover, we will employ the finite-size scaling technique with logarithmic correction proposed by Y.D. Hsieh et al.Hsieh et al. (2013). We also calculate the vortex density and the energy formation of vortex pairs, which may provide insight into the stability of the honeycomb lattice compared to the square lattice and the location of the critical region.
@@ II The model
The XY model represents a system of spins arranged at the sites of a two-dimensional latticeKosterlitz and Thouless (1973) and serves as a powerful tool for modeling magnetic systems in two dimensions, such as magnetic crystalsGong and Zhang (2019) and van der Waals heterostructuresGong et al. (2017).
@@
However, in the XY model, there is no discontinuity in the first derivatives of the free energy at the transition temperature, but rather in the second derivatives. Furthermore, there is no spontaneous symmetry breaking of the system, but rather a change in the nature of quasi-long-range fluctuations of the order parameter, i.e., there are no abrupt changes of the system during the transition. This configures a phase transition of continuous (infinite) orderBerezinskii (1972); Kosterlitz (1974); Kosterlitz and Thouless (1973). For sufficiently low temperatures, there is the creation of vortex-antivortex pairs in the lattice, representing a state with almost long-range order, as predicted by Kosterlitz and ThoulessKosterlitz and Thouless (1973). Vortices can be characterized as topological defects, where the spins arrange themselves in a whirlpool-like formation, that is, when adding the phase differences between the spins in a closed path proportional to the lattice geometry, it yields a value multiple of ±$\pm 2\pi$. As the temperature increases, a transition from this state to a disordered one occurs, breaking the vortex-antivortex pairs, which begin to move freely through the system.
@@ III.1 The Metropolis Algorithm
The MC method uses random numbers to construct the phase space of a given physical system, which can be studied by sampling its microstates. This capability of the MC method is convenient when one wants to represent complex and random systems with fair precisionLandau and Binder (2014).
@@
By scanning the entire lattice, that is, all $N$ sites, where $N=L^{2}$, assigning a new orientation to the spin at each site and using the Metropolis algorithm to accept or not the new orientation, we perform what is known as the MC sweep ($m_{s}$).
@@ III.2 Simulated Annealing
Due to the complexity of the phase space of the XY model, the Metropolis algorithm might face challenges in sampling the low-temperature region. An effective technique to address this issue is the Simulated AnnealingLandau and Binder (2014). This technique consists of initializing the system at an infinite temperature and cooling it uniformly and slowly until $T\approx 0$. This protocol is efficient as it enables the system not to get stuck in local energy minima when minimizing a function dependent on numerous parameters. Therefore, the SA procedure allows the system to pass through local minima without getting stuck in them, as long as the increase in $\beta=1/T$ is small in each interval.
In practice, we start with $\beta=0$ and increase it uniformly until we reach a desired temperature value. Each simulation consists of $n$ temperature intervals, where within each interval $m_{s}$ MC sweeps are performed, resulting in a total of $nm_{s}$ simulation steps. After each MC step, the increments in $\beta$ are given by $\beta=1/k_{B}T=\beta_{0}+i\Delta\beta_{n}\quad(i=1,...,nm_{s})$, where $\Delta\beta_{n}=(\beta_{n}-\beta_{0})/nm_{s}$.
@@ III.3 Wang-Landau simulation
In the study of phase transitions in physical systems, entropic simulations have proven to be an extremely useful toolCaparica and DaSilva (2015); Caparica and Cunha-Netto (2012). Based on the Wang-Landau scheme, they are performed by a random walk in energy space, driven by an unbiased change in spin states, to construct the density of states $g(E)$. The probability of accepting a new spin state is proportional to the reciprocal of the density of states, $1/g(E)$. During the simulation run, a histogram of energy values, $H(E)$, is accumulated, and the usual flatness criterion for $H(E)$ is set $90\%$. After estimating the density of states, these simulations allow the calculation of average values of relevant thermodynamic quantities at any temperature, significantly expanding the information available about the system. In our work, we use entropic sampling based on the WL methodWang and Landau (2001a, b), with some improvements, such as the adoption of the MC sweep to update the density of states, the accumulation of microcanonical average is done only after a more advanced WL level, the use of an automatic criterion to end the simulation run, and the use of an improvement that saves computational time, which consists of starting all the rounds from a single previous simulation up to some WL level and performing all the other runs from thereCaparica and Cunha-Netto (2012); Caparica (2014); Caparica and DaSilva (2015).
@@ III.4 Thermodynamic quantities
To characterize the thermodynamic behavior of the XY model, we calculated the average energy per site, magnetization, and the associated thermal fluctuations, including specific heat and susceptibility. We also analyzed the behavior of the fourth-order cumulant of the order parameter. The average energy in this case is given by:
@@
where $N$ is the total number of spins, $\cal H$ is the Hamiltonian of the system given by Eq. 1, and its average is taken at a fixed temperature.
Although magnetization is not the most suitable order parameter for determining the BKT transition temperature, it plays a significant role in qualitatively understanding the phase transition. Therefore, the magnetization of the XY model on a finite lattice can be expressed as:
@@
From Eq. 7 it can be inferred that $m=1$ corresponds to the ground state and is expected to decrease as the temperature rises. For the specific heat and susceptibility we have,
@@
It is suggested that the second-order helicity modulus serves as an appropriate order parameter to investigate the BKT transitionMinnhagen and Kim (2003); Hsieh et al. (2013). This quantity represents the response of the system to a global twist applied to the spins along a specific direction. To better conceptualize helicity, consider a square lattice of spins at a temperature $T=0$ in a ferromagnetic phase, with linear size $L$, and subjected to a magnetic field parallel to one of its edges. The spins located at the edge that is under the action of the magnetic field must undergo a change of $\Phi$ in their orientation relative to the original state. To minimize the system energy, each spin of neighboring rows parallel to the edge will undergo a rotation equivalent to an angle of $\phi=\Phi/(L-1)$. Thus, the energy in the presence of this spin is given by $E(\phi)=-J(L-1)L\cos(\phi)=E(0)+J(L-1)L[1-\cos(\phi)]$. According to Sandvik et al. (2010), for a small $\phi$, we have $E(\phi)-E(0)=(J/2)(L-1)L\phi^{2}$. Using this result, one can see that the change in the energy for a small twist corresponds to a function of $\phi^{2}$, thus, the helicity for $T=0$ is defined as
@@
because the odd terms of the expansion are null. Thus, the helicity and fourth-order helicity modulus are defined as follows:
@@
As described in Minnhagen and Kim (2003), for a potential dependent solely on the angular difference between spins, the helicity modulus is expressed as:
@@
$s=\frac{1}{N}\sum_{\langle ij\rangle_{x}}\sin(\theta_{i}-\theta_{j}),$ (19)
@@
Therefore, $\langle\Upsilon\rangle\geq 0$. For small variations of $\phi$, the free energy is dominated by the smallest non-zero derivative, which is positive. Moreover, $\langle\Upsilon_{4}\rangle$ must remain positive for any temperature $T$ where $\langle\Upsilon\rangle=0$. This implies that $\langle\Upsilon\rangle$ cannot continually go to zero at the transition temperature $T_{BKT}$ if $\langle\Upsilon_{4}\rangle$ at the same time approaches a minimum at $T_{BKT}$. Consequently, if $\langle\Upsilon_{4}\rangle$ approaches the minimum in $T_{BKT}$, then the change in $\langle\Upsilon\rangle$ must be discontinuousMinnhagen and Kim (2003). Thus, $\langle\Upsilon\rangle$ must exhibit a discontinuity, supporting Minnhagen’s conclusion that both helicity orders can serve as order parameters for the BKT transitionMinnhagen and Kim (2003).
@@
This equality holds in the thermodynamic limit, as the system size draw near to infinity. The procedure involves calculating the helicity modulus for each temperature across various lattice sizes and comparing the results with the curve $r=2/(\beta\pi)$. The intersection points determine the critical temperature for each lattice size. To estimate the BKT transition temperature using the fourth-order helicity modulus (Eq. 20), one simply identifies the temperature at which the function reaches its minimum.
@@
To address the limitation imposed by the finite size of the system in the simulation, we will employ the Finite-Size Scaling (FSS) technique, as developed by K. BinderBinder (1981). For temperatures below the critical value, some physical quantities of interest, such as helicity $\Upsilon$, can be described by a power lawSilva et al. (2022), as $\Upsilon\propto L^{-x}$, where $x$ is an exponent dependent on the system temperature. When the order parameter has a strong dependence on the size of the system and it is possible to find a singularity in the phase transition, a linear extrapolation ($x=1$) is appropriate to find the critical temperature in the thermodynamic limit ($L\rightarrow\infty$). Otherwise, it cannot be guaranteed that the order parameter is discontinuous in this limit. This appears to be the case for second-order helicity. However, to mitigate this challenge, some authors propose the use of a logarithmic correction, as described inLandau and Binder (2014); Silva et al. (2022); Binder (1997):